OpenStudy (harvey03):

Some material has a density of 5g/cm3. Calculate the mass of 30cm3 of this material and give the unit.

OpenStudy (rushwr):

Density = mass divided by volume \[density = \frac{ mass }{ volume }\] when u subject mass : \[mass = density * volume \]

OpenStudy (rushwr):

mass = 5 * 30 = 150g

OpenStudy (harvey03):

Thanks!

OpenStudy (rushwr):

no problem : @harvey03